# Inhomogeneous SSH models and the doubling of orthogonal polynomials

### v1

@@ 3 Inhomogeneous SSH model from the doubling procedure
$Q_{2n+1}(x)=\frac{\tau_{2}t_{n}^{-}x}{(\pi_{x}-\tau_{0})}\left(V_{n+1}(\pi_{x})+\frac{t_{n}^{+}}{t_{n}^{-}}V_{n}(\pi_{x})\right)\,.$ (3.19)
@@
$\frac{\tau_{2}t_{n}^{-}x}{(\pi_{x}-\tau_{0})}\left(t_{n+1}^{+}R_{n+1}(\pi_{x})+\left(t_{n}^{-}+\frac{(t_{n}^{+})^{2}}{t_{n}^{-}}\right)R_{n}(\pi_{x})+\frac{t_{n}^{+}}{t_{n}^{-}}t_{n-1}^{-}R_{n-1}(\pi_{x})\right)$
$=\frac{\tau_{2}t_{n}^{-}x}{(\pi_{x}-\tau_{0})}\left(\frac{\epsilon\sqrt{A_{n}C_{n+1}}}{\tau_{2}t_{n}^{-}}R_{n+1}(\pi_{x})-\frac{(A_{n}+C_{n}+\tau_{0})}{\tau_{2}t_{n}^{-}}R_{n}(\pi_{x})+\frac{\epsilon\sqrt{A_{n-1}C_{n}}}{\tau_{2}t_{n}^{-}}R_{n+1}(\pi_{x})\right)$
$=\frac{\tau_{2}t_{n}^{-}x}{(\pi_{x}-\tau_{0})}\frac{(\pi_{x}-\tau_{0})R_{n}(\pi_{x})}{\tau_{2}t_{n}^{-}}=Q_{2n+1}(x),$ (3.20)
@@
$=\frac{1}{\tau_{2}}(\epsilon\sqrt{A_{n}C_{n+1}}R_{n+1}(\tau_{0})-(A_{n}+C_{n}+\tau_{0})R_{n}(\tau_{0})+\epsilon\sqrt{A_{n-1}C_{n}}R_{n-1}(\tau_{0}))$
@@
$Q_{2n+1}(x)=\frac{\tau_{2}t_{n}^{-}x}{(\pi_{x}-\tau_{0})}\left(V_{n+1}(\pi_{x})-\frac{V_{n+1}(\tau_{0})}{V_{n}(\tau_{0})}V_{n}(\pi_{x})\right)\,$, (3.22)
@@
In the following, we use the above construction with well-known finite sequences of orthogonal polynomials from the ($q$-)Askey scheme (see e.g. [32]) to provide examples of inhomogeneous generalizations of the SSH model, and hence construct interesting solvable physical models.
@@ 4 Inhomogeneous SSH model associated to Krawtchouk polynomials
In this section, we give a solution to the constraints of Proposition 3.1 associated to the Krawtchouk polynomials. These polynomials are defined as follows in terms of the $_{2}F_{1}$ hypergeometric function, for $0<p<1$ and $0\leqslant n\leqslant N-1$:
$K_{n}(x;p,N-1)={}_{2}F_{1}\left({{-n,\;-x}\atop{-N+1}}\;\Bigg|\frac{1}{p}\right)\,$, (4.1)
@@
The grid associated to this sequence of polynomials is $\lambda(k)=-k$, ($k=0,1,\dots,N-1$). The constraints (3.9b) and (3.9c) read
@@
A solution of these constraints is
@@
With these choices, Proposition 3.1 allows us to diagonalize exactly the following $(2N+1)\times(2N+1)$ matrix,
@@
Using Proposition 3.1 and after manipulating the Krawtchouk polynomials, the components of the eigenvectors $\mathbb{Q}(x_{i}^{\pm})$ ($i=0,1,\dots,N-1$) are
@@
To prove relation (4.11a), we used the dual of the backward shift relation of the Krawtchouk polynomials [32] (or the contiguity relation $(KI)$ of Ref. [34]), i.e.,
@@
The components of the eigenvector $\mathbb{Q}(0)$, associated to the vanishing eigenvalue, are:
$Q_{2n}(0)=(-1)^{n}\sqrt{\frac{Np^{n+1}}{(1-p)^{n}}\binom{N}{n}}\,,\quad Q_{2n+1}(0)=0\,.$ (4.13)
The eigenvectors $\mathbb{Q}(x_{k}^{\pm})$, $\mathbb{Q}(0)$ are orthogonal, with norms
$\mathbb{Q}(x_{k}^{\pm})^{t}\ \mathbb{Q}(x_{k}^{\pm})=\frac{2(k+1)(1-p)^{i-N+1}}{p^{i}\binom{N-1}{i}}\,,\qquad\mathbb{Q}(0)^{t}\ \mathbb{Q}(0)=\frac{pN}{(1-p)^{N}}\,$, (4.14)
@@ 5 Inhomogeneous SSH model associated to $q$-Racah polynomials
In the case of the Krawtchouk polynomials discussed in the previous section, a significant simplification arises in the computation (4.11a) of $Q_{2n}(x)$, since it can be expressed as a single Krawtchouk polynomial. The relations of type (4.12) used to achieve this simplification are called contiguity relations, and analogous relations exist for other finite sequences in the ($q$-)Askey scheme. This construction is now extended to the $q$-Racah polynomials, making use of the classification of contiguity relations provided in [34].
@@ 5.1 Properties of the $q$-Racah polynomials
The $q$-Racah polynomials $P^{(qR)}_{n}$ are defined in terms of $q$-hypergeometric functions as follows. For $N$ a nonnegative integer and $0\leqslant n\leqslant N$,
$P^{(qR)}_{n}(\lambda_{x};{\bm{\rho}})={}_{4}\phi_{3}\left({{q^{-n},\;\alpha\beta q^{n+1},\;q^{-x},\;\gamma\delta q^{x+1}}\atop{\alpha q,\;\beta\delta q,\;\gamma q}}\;\Bigg|\;q;q\right)\,$, (5.1)
@@
$P_{n}^{(qR)}(\lambda_{x};{\bm{\rho}})=P_{x}^{(qR)}(\lambda_{n}^{d};{\bm{\rho}}^{d})$ (5.8)
for $0\leqslant n,x\leqslant N$, where $\lambda_{n}^{d}=q^{-n}+\alpha\beta q^{n+1}$ and ${\bm{\rho}}^{d}=\gamma,\delta,\alpha,\beta.$ Furthermore, the following identity holds
@@ 5.2 First exactly solvable model
With these solutions, the results of Proposition 3.1 allow one to exactly diagonalize the corresponding $(2N+1)\times(2N+1)$ matrix given in (3.7), whose spectrum is
@@
By applying Proposition 3.1 and performing suitable manipulations of the $q$-Racah polynomials, the components of the eigenvectors $\bm{Q}(x_{k}^{\pm})$, for $k=0,1,\dots,N-1$, are given by
$Q_{2n}(x_{k}^{\pm})=\epsilon^{n}(-1)^{\nu_{n}}\sqrt{\frac{(1-\beta q)(\delta-\alpha q)(1-q^{-N})w(n,\overline{{\bm{\rho}}}^{d})}{(1-\alpha\beta q^{2})}}P_{n}^{(qR)}(\lambda_{k};\overline{{\bm{\rho}}})\,,\quad(n=0,\ldots,N)\,$, (5.15a)
$Q_{2n+1}(x_{k}^{\pm})=\pm\epsilon^{n}(-1)^{\nu_{n}}\sqrt{(1-q^{k-N})(\delta-q^{-k})w(n,{\bm{\rho}}^{d})}P_{n}^{(qR)}(\lambda_{k};{\bm{\rho}})\,,\quad(n=0,\ldots,N-1)\,.$ (5.15b)
@@
The components of the eigenvector $\mathbb{Q}(0)$, corresponding to the vanishing eigenvalue, are given by
$Q_{2n}(0)=\epsilon^{n}(-1)^{\nu_{n}}\sqrt{\frac{(1-\beta q)(\delta-\alpha q)(1-q^{-N})w(n,\overline{{\bm{\rho}}}^{d})}{(1-\alpha\beta q^{2})}}P_{n}^{(qR)}(\lambda_{N};\overline{{\bm{\rho}}})\,,\quad Q_{2n+1}(0)=0\,.$ (5.18)
The eigenvectors $\mathbb{Q}(x_{k}^{\pm})$, $\mathbb{Q}(0)$ are orthogonal, with norms, for $k=0,1,\dots,N-1$:
@@
$\qquad\mathbb{Q}(0)^{t}\ \mathbb{Q}(0)$ $=\frac{(1-\beta q)(\delta-\alpha q)(1-q^{-N})}{(1-\alpha\beta q^{2})}h_{N,\overline{{\bm{\rho}}}^{d}}\,.$ (5.20)
@@ 5.3 Second exactly solvable model
Additional contiguity relations for $q$-Racah polynomials were derived in [34]. Some of these relations do not alter the value of $N$, in contrast to those used previously (see (4.12) or (5.17)). By employing such relations, alternative solutions to the constraints can be obtained but they apply to a chain with an even number of sites and a slight modification of Proposition 3.1 is needed.
@@
In this case, $t_{N-1}^{-}=0$. We can eliminate the last row and column of the $(2N+1)\times(2N+1)$ matrix and consider the resulting subsystem. By adapting the proof of Proposition 3.1, we can show that the spectrum of the corresponding $2N\times 2N$ matrix is given by
@@
$Q_{2n}(x_{k}^{\pm})=(-1)^{n}\sqrt{\frac{(1-\beta q)(1-\alpha\beta q^{N+1})(1-\beta\delta q)w(n;\overline{{\bm{\rho}}}^{d})}{1-\alpha\beta q^{2}}}P_{n}^{(qR)}(\lambda_{k};\overline{{\bm{\rho}}}),$ (5.26a)
$Q_{2n+1}(x_{k}^{\pm})=\pm(-1)^{n}\sqrt{(1-\delta\beta q^{k+1})(1-\beta q^{N-k})w(n;{\bm{\rho}}^{d})}P_{n}^{(qR)}(\lambda_{k};{\bm{\rho}}),$ (5.26b)
for $n=0,\ldots,N-1$. The parameters are specified as follows:
@@
We note that there is no shift in the parameter N1N-1 when comparing $\bm{\rho}$} and $\overline{{\bm{\rho}}}$. The eigenvectors $\mathbb{Q}(x_{i}^{\pm})$ are orthogonal, with norms, for $k=0,1,\dots,N-1$:
$\mathbb{Q}(x_{k}^{\pm})^{t}\ \mathbb{Q}(x_{k}^{\pm})=2(1-\delta\beta q^{k+1})(1-\beta q^{N-k})h_{k,{\bm{\rho}}^{d}}\,.$ (5.29)
@@ 6 Conclusion
A natural next step is to investigate how the topological properties of the SSH model, along with other physically relevant quantities such as entanglement measures and correlation functions, are influenced by the introduction of these inhomogeneities. Since the doubling procedure provides a clear analytical handle on both the energy spectrum and the eigenstates, we expect that the computation and interpretation of these more intricate quantities will remain accessible through analytical methods. We will explore these aspects in detail in a forthcoming paper.

### v2

@@ 3.4 Connection with the doubling procedure
$Q_{2n+1}(x)=\frac{\tau_{2}t_{n}^{-}x}{\pi_{x}-\tau_{0}}\left(V_{n+1}(\pi_{x})+\frac{t_{n}^{+}}{t_{n}^{-}}V_{n}(\pi_{x})\right)\,.$ (3.23)
@@
$\frac{\tau_{2}t_{n}^{-}x}{\pi_{x}-\tau_{0}}\left(t_{n+1}^{+}R_{n+1}(\pi_{x})+\left(t_{n}^{-}+\frac{(t_{n}^{+})^{2}}{t_{n}^{-}}\right)R_{n}(\pi_{x})+\frac{t_{n}^{+}}{t_{n}^{-}}t_{n-1}^{-}R_{n-1}(\pi_{x})\right)$
$=\frac{\tau_{2}t_{n}^{-}x}{\pi_{x}-\tau_{0}}\left(\frac{\epsilon\sqrt{A_{n}C_{n+1}}}{\tau_{2}t_{n}^{-}}R_{n+1}(\pi_{x})-\frac{(A_{n}+C_{n}+\tau_{0})}{\tau_{2}t_{n}^{-}}R_{n}(\pi_{x})+\frac{\epsilon\sqrt{A_{n-1}C_{n}}}{\tau_{2}t_{n}^{-}}R_{n+1}(\pi_{x})\right)$
$=\frac{\tau_{2}t_{n}^{-}x}{\pi_{x}-\tau_{0}}\ \frac{\pi_{x}-\tau_{0}}{\tau_{2}t_{n}^{-}}R_{n}(\pi_{x})=Q_{2n+1}(x),$ (3.24)
@@
$=\frac{1}{\tau_{2}}\left(\epsilon\sqrt{A_{n}C_{n+1}}R_{n+1}(\tau_{0})-(A_{n}+C_{n}+\tau_{0})R_{n}(\tau_{0})+\epsilon\sqrt{A_{n-1}C_{n}}R_{n-1}(\tau_{0})\right)$
@@
$Q_{2n+1}(x)=\frac{\tau_{2}t_{n}^{-}x}{\pi_{x}-\tau_{0}}\left(V_{n+1}(\pi_{x})-\frac{V_{n+1}(\tau_{0})}{V_{n}(\tau_{0})}V_{n}(\pi_{x})\right)\,$, (3.26)
@@
In the following, we use the above construction with well-known finite sequences of orthogonal polynomials from the ($q$-)Askey scheme (see e.g. [34]) to provide examples of inhomogeneous generalizations of the SSH model, and hence construct interesting solvable physical models. The results for each family of polynomials discussed in this work, including their corresponding couplings and eigenvalues, are summarized in the following table:
Polynomials $t^{+}_{n}$ $t^{-}_{n}$ $\tau_{2}$ $\tau_{0}$ Eigenvalues
[…]
Krawtchouk $\sqrt{p(N-n)}$ $\sqrt{(1-p)(n+1)}$ 1-1 11 0,±k+10,\ $\pm\sqrt${k+1}
[…]
Table 1: Summary of the different SSH models studied in this paper.
@@ 4 Inhomogeneous SSH model associated to Krawtchouk polynomials
4.1 Properties of Krawtchouk polynomials
In this section, we give a solution to the constraints (3.11) associated to the Krawtchouk polynomials. These polynomials are defined as follows in terms of the $_{2}F_{1}$ hypergeometric function:
$K_{n}(x;p,N-1)={}_{2}F_{1}\left({{-n,\;-x}\atop{-N+1}}\;\Bigg|\frac{1}{p}\right)\,,\quad(0<p<1\,,\quad n=0,1,\dots,N-1)\,$, (4.1)
@@ 4.1 Properties of Krawtchouk polynomials
The grid associated to this sequence of polynomials is $\lambda(k)=-k$, ($k=0,1,\dots,N-1$).
4.2 SSH model of Krawtchouk type
For this choice of polynomials, the constraints (3.11b) and (3.11c) read
@@ 4.2 SSH model of Krawtchouk type
A solution to these constraints is
@@
With these choices, results of Sec. 3 allow us to diagonalize exactly the following $(2N+1)\times(2N+1)$ matrix,
@@
Using results of Sec. 3 and after manipulating the Krawtchouk polynomials, the components of the eigenvectors $\mathbb{Q}(x_{i}^{\pm})$ ($i=0,1,\dots,N-1$) are
@@
To prove relation (4.11a), we used the contiguity relation $(KI)$ of Ref. [41] (also known as the backward shift relation of the Krawtchouk polynomials [34] up to self duality), i.e.,
@@
The components of the eigenvector $\mathbb{Q}^{(0)}$, associated to the vanishing eigenvalue (or zero mode), are:
$Q_{2n}^{(0)}=(-1)^{n}\sqrt{\frac{p^{n}}{(1-p)^{n}}\binom{N}{n}}\,,\quad Q_{2n+1}^{(0)}=0\,.$ (4.13)
The zero mode amplitude $\left(Q_{2n}^{(0)}\right)^{2}$ is maximal close to $n\sim pN$, which corresponds to the position in the chain where $t_{n}^{+}\sim t_{n}^{-}\sim\sqrt{Np(1-p)}$ for the Krawtchouk case. Because $t_{n}^{+}$ (resp. $t_{n}^{-}$) is monotonically decreasing (increasing) with $n$, the zero mode indeed separates regions where $t_{n}^{+}>t_{n}^{-}$ and $t_{n}^{+}<t_{n}^{-}$, in agreement with our general discussion of Sec. 3.
The eigenvectors $\mathbb{Q}(x_{k}^{\pm})$, $\mathbb{Q}^{(0)}$ are orthogonal, with norms
$\mathbb{Q}(x_{k}^{\pm})^{t}\ \mathbb{Q}(x_{k}^{\pm})=\frac{2(k+1)(1-p)^{i-N+1}}{p^{i}\binom{N-1}{i}}\,,\qquad{\mathbb{Q}^{(0)}}^{t}\ \mathbb{Q}^{(0)}=\frac{1}{(1-p)^{N}}\,$, (4.14)
@@ 5 Inhomogeneous SSH model associated to $q$-Racah polynomials
In the case of the Krawtchouk polynomials discussed in the previous section, a significant simplification arises in the computation (4.11a) of $Q_{2n}(x)$, since it can be expressed as a single Krawtchouk polynomial. The relations of type (4.12) used to achieve this simplification are called contiguity relations, and analogous relations exist for other finite sequences in the ($q$-)Askey scheme. This construction is now extended to the $q$-Racah polynomials, making use of the classification of contiguity relations provided in [41]. In particular, in the following we define two different solvable models associated to these polynomials.
@@ 5.1 Properties of the $q$-Racah polynomials
The $q$-Racah polynomials $P^{(qR)}_{n}$ are defined in terms of $q$-hypergeometric functions as follows, for $N$ a nonnegative integer
$P^{(qR)}_{n}(\lambda_{x};{\bm{\rho}})={}_{4}\phi_{3}\left({{q^{-n},\;\alpha\beta q^{n+1},\;q^{-x},\;\gamma\delta q^{x+1}}\atop{\alpha q,\;\beta\delta q,\;\gamma q}}\;\Bigg|\;q;q\right)\,,\quad(n=0,1,\dots,N)\,$, (5.1)
@@
$P_{n}^{(qR)}(\lambda_{x};{\bm{\rho}})=P_{x}^{(qR)}(\lambda_{n}^{d};{\bm{\rho}}^{d})\,,\quad(n,x=0,1,\dots N)\,$, (5.8)
where $\lambda_{n}^{d}=q^{-n}+\alpha\beta q^{n+1}$ and ${\bm{\rho}}^{d}=\gamma,\delta,\alpha,\beta.$ Furthermore, the following identity holds
@@ 5.2 First exactly solvable model
With these solutions, the general results obtained in Sec. 3 allow one to exactly diagonalize the corresponding $(2N+1)\times(2N+1)$ matrix given in (3.12), whose spectrum is
@@
Using results of Sec. 3 and performing suitable manipulations of the $q$-Racah polynomials, the components of the eigenvectors $\bm{Q}(x_{k}^{\pm})$, for $k=0,1,\dots,N-1$, are given by
$Q_{2n}(x_{k}^{\pm})=\epsilon^{n}(-1)^{\nu_{n}}\sqrt{\frac{(1-\beta q)(\delta-\alpha q)(1-q^{-N})w(n,\overline{{\bm{\rho}}}^{d})}{1-\alpha\beta q^{2}}}P_{n}^{(qR)}(\lambda_{k};\overline{{\bm{\rho}}})\,$, $(n=0,\ldots,N)\,$, (5.15a)
$Q_{2n+1}(x_{k}^{\pm})=\pm\epsilon^{n}(-1)^{\nu_{n}}\sqrt{(1-q^{k-N})(\delta-q^{-k})w(n,{\bm{\rho}}^{d})}P_{n}^{(qR)}(\lambda_{k};{\bm{\rho}})\,$, $(n=0,\ldots,N-1)\,.$ (5.15b)
@@
The components of the eigenvector $\mathbb{Q}^{(0)}$, corresponding to the vanishing eigenvalue, are given by
$Q_{2n}^{(0)}=(-1)^{n}\sqrt{\frac{(q^{N}\delta)^{n}(1-\alpha\beta q^{2n+1})(q^{-N},\alpha\beta q,\beta q,\alpha q/\delta;q)_{n}}{(1-\alpha\beta q)(q,\beta\delta q,\alpha\beta q^{N+2},\alpha q;q)_{n}}}\,,\quad Q_{2n+1}^{(0)}=0\,.$ (5.18)
The eigenvectors $\mathbb{Q}(x_{k}^{\pm})$, $\mathbb{Q}^{(0)}$ are orthogonal, with norms, for $k=0,1,\dots,N-1$:
@@
$\qquad{\mathbb{Q}^{(0)}}^{t}\ \mathbb{Q}^{(0)}$ $=\frac{(\delta,\alpha\beta q^{2};q)_{N}}{(\alpha q,\beta\delta q;q)_{N}}\,.$ (5.20)
@@ 5.3 Second exactly solvable model
Additional contiguity relations for $q$-Racah polynomials were derived in [41]. Some of these relations do not alter the value of $N$, in contrast to those used previously (see (4.12) or (5.17)). By employing such relations, alternative solutions to the constraints can be obtained but they apply to a chain with an even number of sites and a slight modification of results of Sec. 3 is needed.
@@
In this case, $t_{N-1}^{-}=0$. We can eliminate the last row and column of the $(2N+1)\times(2N+1)$ matrix and consider the resulting subsystem. By adapting the proof given in Sec. 3, we can show that the spectrum of the corresponding $2N\times 2N$ matrix is given by
@@
$Q_{2n}(x_{k}^{\pm})=(-1)^{n}\sqrt{\frac{(1-\beta q)(1-\alpha\beta q^{N+1})(1-\beta\delta q)w(n;\overline{{\bm{\rho}}}^{d})}{1-\alpha\beta q^{2}}}P_{n}^{(qR)}(\lambda_{k};\overline{{\bm{\rho}}}),$ $(n=0,\ldots,N-1)$ (5.26a)
$Q_{2n+1}(x_{k}^{\pm})=\pm(-1)^{n}\sqrt{(1-\delta\beta q^{k+1})(1-\beta q^{N-k})w(n;{\bm{\rho}}^{d})}P_{n}^{(qR)}(\lambda_{k};{\bm{\rho}}),$ $(n=0,\ldots,N-1)\,.$ (5.26b)
The parameters are specified as follows:
@@
We note that there is no shift in the parameter N1N-1 when comparing $\bm{\rho}$} and $\overline{{\bm{\rho}}}$. The eigenvectors $\mathbb{Q}(x_{i}^{\pm})$ are orthogonal, with norms:
$\mathbb{Q}(x_{k}^{\pm})^{t}\ \mathbb{Q}(x_{k}^{\pm})=2(1-\delta\beta q^{k+1})(1-\beta q^{N-k})h_{k,{\bm{\rho}}^{d}}\,,\quad(k=0,1,\dots,N-1)\,.$ (5.29)
@@ 6 Conclusion
A natural next step is to investigate how the topological properties of the SSH model, along with other physically relevant quantities such as entanglement measures and correlation functions, are influenced by the introduction of these inhomogeneities. Since the doubling procedure provides a clear analytical handle on both the energy spectrum and the eigenstates, we expect that the computation and interpretation of these more intricate quantities will remain accessible through analytical methods. We will explore these aspects in detail in a forthcoming paper [38].
Finally, it would be insightful to determine whether other physically relevant models can be analyzed using the approach proposed in this paper. For instance, the Kitaev chain with alternating bulk chemical potential, as explored in [42], could be generalized and solved exactly. Furthermore, the doubling of bivariate orthogonal polynomials—a topic not yet studied in detail—could facilitate the investigation of higher-dimensional inhomogeneous topological models. This would allow for a direct comparison with previous exact results for the $n$-dimensional SSH model [43]. Additionally, extending this study to the inhomogeneous SSH model with periodic boundary conditions would be of great interest; however, the theory of the required special functions is currently less developed and necessitates further mathematical research.